OpenStudy (anonymous):

The formula for volume of this rectangular prism is: V = 2x 3 + 17x 2 + 46x + 40 Find an expression for the missing side length.

OpenStudy (anonymous):

use mathway.com it has saved me many times.

OpenStudy (anonymous):

okay but I need to know how to solve it

OpenStudy (anonymous):

Oh okay well that I can't help you with but if you desperately need an answer that will work.

OpenStudy (anonymous):

thanks

OpenStudy (the_fizicx99):

(x + 2)(x+4) = x^2 + 4x + 2x + 8 = x^2 + 8x + 8

OpenStudy (the_fizicx99):

V = LWH

OpenStudy (the_fizicx99):

Oh wait! no i messed up X_X

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

its fine you are still helping some

OpenStudy (the_fizicx99):

You're suppose to divide that, no i didn't mess up xDD 2x^3 + 17x^2 + 46x + 40/ x^2 + 8x + 8 Thats how you'll find the missing side length

OpenStudy (anonymous):

THANK YOU

OpenStudy (the_fizicx99):

Shiet 4x + 2x = 6x not 8x 2x^3 + 17x^2 + 46x + 40/ x^2 + 6x + 8 ... Sorry, i'm in a rush D:

OpenStudy (anonymous):

no lol thank you

OpenStudy (the_fizicx99):

yw